# Graph-Informed Neural Networks for Sparse Grid-Based Discontinuity Detectors

### v1

@@ 5.1 Two-dimensional Discontinuity Detection
We start the numerical experiments testing our method on 2-dimensional test functions; we will use the notation introduced after (4), p.4, for describing the sparse grids.
Let $\mathcal{S}_{2}$ be the 2-dimensional sparse grid $\mathcal{S}_{2}:=\mathcal{I}_{\rm sum}(6)$ ($N=65$ points, see Figures 1 and 2). Then, we randomly generate a set $\mathcal{G}_{2}$ of $Q=600$ piece-wise continuous functions characterized by a linear, spherical, and polynomial cut identified by the zero-level set of given functions (see Section A.2 for details); specifically, we have 200 functions for each cut type. Then, we build the synthetic dataset of pairs $(\boldsymbol{g}^{\prime},\boldsymbol{p})$ from $\mathcal{G}_{2}$, running Algorithm 1 and detecting the zero-level sets containing the discontinuity interfaces of the functions in $\mathcal{G}_{2}$ (we use detector $Z^{(150)}$, see Section A.1 for details).
@@
  • Loss function: loss (19), with $w_{0}=0.5$ and $w_{1}=1.5$;
@@
Concerning the NN-based detectors, we use a leaky relu activation function [1], a Glorot Normal initialization for the weights [11], and (for the GINN-based detector) $F=15$ filters for the hidden layers. At the end of the training, we measure the loss and, for a better understanding, the Mean Absolute Error (MAE) of these models on the test set $\mathcal{P}_{2}$ (see Table 1). Looking at the values in the table, we see that the GINN's prediction performances are slightly better than the MLP's ones.
Loss (see (19)) MAE
MLP 4.8678 0.0480
GINN 3.5396 0.0393
@@
Now, we test the discontinuity detection abilities of these NN models if used as discontinuity detectors together with the sparse grid-based algorithm proposed (see Algorithm 2). In particular, we test them on four piece-wise continuous functions defined on $\Omega_{2}=[-1,1]^{2}$ (see Figure 5): $i$) a function with circular cut; $ii$) a function with Legendre polynomial cut; $iii$) a function with sinusoidal cut; $iv$) a function with one elliptic cut and two bow-shaped cuts. We point the attention of the reader to the fact that functions ($i$) and ($ii$) are of the same type of functions used to generate the synthetic dataset $\mathcal{D}_{2}$. On the other hand, functions ($iii$) and ($iv$) are functions of a different type with respect to the ones in $\mathcal{G}_{2}$; then, they are somehow ``new'' to the NN models and good performances on these functions mean good generalization abilities of the NNs. Moreover, looking at Figure 5, we can observe that test functions ($i$), ($ii$), and ($iv$) have discontinuity interfaces characterized by parts where the discontinuity jump are approximately zero, i.e., where the interface is not of (evident) discontinuity; then, they are important to understand how much sensitive the NN-based detectors are in identifying discontinuities.
@@
The performances of the NN-based detectors are measured according to a True Positive Rate (TPR). Let $T$ be the set of final troubled points detected by the method and let $\Lambda_{\min}$ be the minimum edge length used as stopping criterion (see Algorithms 1 and 2); then, the TPR with respect to each test functions is computed in this way:
@@
    let $\zeta$ be the zero-level set of the function characterizing the discontinuity interface of the test function. Let $\widetilde{\mathcal{S}}_{2}$ be the 2-dimensional sparse grid $\mathcal{I}_{\rm sum}(3)$ built with respect to the collocation knots (3);
@@
For measure the detection quality of our method we focus on the TPR described above. We do not consider a True Negative Rate (TNR) based on the same principles because it would be uninformative, even in the case of exact detectors. In fact, given the non-troubled points returned by Algorithm 1 using an exact detector $\Delta^{*}$ (i.e., all the points evaluated during the procedure and not belonging to $T$), the discontinuity interface can still intersect one edge of the SGG $\widetilde{G}(\boldsymbol{x})$ for some of these points; the reason is that the non-troubled points have been discarded by the algorithm only because their adjacent points are nearer to the intersection between the discontinuity interface and the graph's edge, not because they are too far from the interface (e.g., see all the non-troubled points of the SGG $\widetilde{G}(\boldsymbol{x})$ in the center of Figure 5(b)).
@@
Figure 6: SGG based on the sparse grid $\widetilde{\mathcal{S}}_{2}$ (left) and its use for computing the TPR (right). Dark green circles are the troubled points detected by the algorithm at the current step. Opaque circles are the troubled points assigned to $T$ (i.e., final troubled points), transparent ones are points assigned to $L$ (i.e., centers of smaller sparse grids in the next algorithm's steps). In this example, the opaque dark green circles are true troubled points because the discontinuity interface intersects at least one edge of the SGGs.
We run Algorithm 2 on the four test functions using the NN-based detectors, and using the same input parameters for all the cases; i.e., same starting set $L$ of center-edge length pairs and same $\Lambda_{\min}=2/2^{(h_{\max}+1)}=2^{-5}$ for stopping criterion ($h_{\max}=5$, maximum discretization level of the sparse grid). Moreover, since during the procedure some sparse grids may have points outside the domain $\Omega_{2}$, we modify Algorithm 2 in order to stop the search if a troubled point is detected outside the domain (i.e., it stops even if the stopping citerion with respect to $\Lambda_{\min}$ is not satisfied).
In Table 2, we report the TPRs, the number of troubled points detected, and the total number of points visited by the two NN-based discontinuity detectors, with respect to the test functions. Looking at the results reported in this table, we observe that both the NN-based detectors have extremely good TPRs (at least $98\%$). According with the test set performances (see Table 1), we see that the GINN-based detector is more precise than the MLP-based one (higher TPRs and TPR equal to $100\%$ in three cases over four); moreover, it is more sensitive for detecting discontinuities (more troubled points detected). In particular, looking at the collocation of the detected troubled points (Figures 7 and 8), we can deduce that the higher sensitiveness of the GINN-based detector depends on its ability of detecting troubled points even if the discontinuity jumps are very small.
@@
NN model Test function TPR Num. of troubled points Total num. of points
@@ 5.1.1 Edge Detection Problem in Images
Black and white images can be modeled as functions $g:\Omega\subset\mathbb{R}^{2}\rightarrow[0,1]$, where $g(\boldsymbol{x})$ is the grayscale level of the pixel, for each pixel $\boldsymbol{x}\in\Omega$. Then, two-dimensional detectors can be used for detecting edges in pictures, since edges are discontinuity interfaces of $g$. Analogously to [30], we test our NN-based two-dimensional detectors on the edge detection problem of the Shepp-Logan phantom [23], varying the image resolution. In particular, let $P=(p_{ij})\in[0,1]^{r\times r}$ be the matrix representing the Shepp-Logan phantom image with resolution $r\times r$ pixels; then, we apply Algorithm 2 modeling the image as a function $g_{r}:\mathbb{R}^{2}\rightarrow[0,1]$ such that:
$g_{r}(\boldsymbol{x})=\begin{cases}p_{[x_{1}]+1,[x_{2}]+1}\,,\quad&\text{if }0\leq[x_{1}],[x_{2}]\leq r-1\\ 0\,,\quad&\text{otherwise}\end{cases}\,,$
where $[x]$ is the integer round-off of $x$, for each $x\in\mathbb{R}$.
In Figure 9, we report the results obtained on $g_{r}$ using Algorithm 2 and the two NN-based discontinuity detectors, for each $r=256,512,1024$. Looking at this figures, we observe that the edge detection performances of the NN-based method increase with the resolution, similarly to the CNN detectors of [30]; however, we see that the GINN-based detector shows very good performances even in the lowest resolution ($256\times 256$), while the MLP-based detector obtains acceptable results only for the highest resolution. Therefore, we have a further experimental proof about the better detection abilities of the GINN-based detector with respect to the MLP-based one.
A very interesting characteristic of our method applied to the edge detection problem can be observed in the rightmost column of Table 3, which reports the number of troubled points detected and the total number of points visited by the NN-based detectors, for each image resolution. Indeed, we observe that increments in the resolution do not imply a considerable higher number of points evaluated by the method. This characteristic is very important because it means that increments in the image resolution guarantee better results at almost the same cost; in particular, for the last two cases, we have that the function evaluations done by the algorithm are (approximately) only $330^{2}$, that is much smaller than the number of pixels available ($512^{2}$ and $1024^{2}$, respectively).
(a) MLP, $256\times 256$
(b) GINN, $256\times 256$
(c) MLP, $512\times 512$
(d) GINN, $512\times 512$
(e) MLP, $1024\times 1024$
(f) GINN, $1024\times 1024$
@@
NN model Image Resolution Num. of troubled points Total num. of points
MLP $256\times 256$ 5222 107201 ($\sim 327^{2}$)
GINN $256\times 256$ 6417 109225 ($\sim 330^{2}$)
@@ 5.2 Four-dimensional Discontinuity Detection
Following the considerations made before, in this subsection we focus on a GINN-based detector, as in general GINNs perform better than MLPs. The sparse grid we consider is the four dimensional generalization of $\mathcal{S}_{2}$; i.e., the sparse grid $\mathcal{S}_{4}:=\mathcal{I}_{\rm sum}(8)$ in $\mathbb{R}^{4}$ ($N=401$ points). Then, we consider a randomly generated set $\mathcal{G}_{4}$ of 750 functions, again with linear, spherical, and polynomial cut (250 each). We use a larger set of functions (with respect to Section 5.1) because we use a less precise approximation of the exact detector for the creation of the synthetic dataset ($Z^{(3)}=\left.Z^{(t+1)}\right|_{t=2}$, see Section A.1 for details); indeed, we prefer a less precise detector to make the dataset creation process less expensive, even if at the cost of a less detailed dataset.
@@
Loss (see (19)) MAE
GINN 81.8518 0.1813
@@
Given the trained GINN-based four-dimensional detector, we test its abilities running Algorithm 2 ($\Lambda_{\min}=2/2^{(h_{\max})}=2^{-4}$) on a piece-wise continuous four-dimensional test function. Specifically, we use a test function $g_{\vartheta}:\Omega_{4}\subseteq\mathbb{R}^{4}\rightarrow\mathbb{R}$ with domain $\Omega_{4}=[-1,1]^{4}$, continuous pieces defined by four-dimensional Legendre polynomials (see (24), Section A.2), and as discontinuity interface a torus with size depending on $x_{4}$; i.e., the discontinuity interface of $g_{\vartheta}$ is characterized by the function
@@
The results of Algorithm 2 for this four-dimensional case are very promising. In Table 5, we report the TPR, the number of troubled points detected, and the total number of points visited by the GINN-based detector; then, looking at this table, we see a TPR of $93.52\%$ for more than half a million of troubled points. Moreover, the total number of points evaluated during the algorithm execution is approximately equal to the number of points of a coarse regular grid of $60$ knots per dimension but the point distribution is more dense near to the discontinuity interface $\vartheta$, instead of being equally distributed in $\Omega_{4}$.
In Figure 10, we report some visual examples of the troubled points identified by the GINN-based detector for the function $g_{\vartheta}$. In particular, for fixed values of $x_{4}$, we show cases where the detector identifies $\vartheta$ almost perfectly with (almost) no false troubled points (Figures 9(a) and 9(b)), with few true troubled points (Figure 9(c)), and with many false troubled points (Figure 9(d)).
NN model TPR Num. of troubled points Total num. of points
@@
Figure 10: Visual examples of the troubled points detected by the GINN-based detector for $g_{\vartheta}$. Each picture is a visualization in $\mathbb{R}^{3}$ of $g_{\vartheta}$ for a fixed value of $x_{4}$. The colored surface is the size-changing torus $\vartheta$ (see (21)). Black dots are true troubled points, red dots are false troubled points.

### v4

@@ 5.1 Two-dimensional Discontinuity Detection
We start the numerical experiments testing our method on 2-dimensional test functions; we will use the notation introduced after equation (2) for describing the sparse grids.
Let $\mathcal{S}_{2}$ be the 2-dimensional sparse grid $\mathcal{S}_{2}:=\mathcal{I}_{\rm sum}(6)$ ($N=65$ points, see Figures 1 and 2). Then, we randomly generate a set $\mathcal{G}_{2}$ of $Q=600$ piece-wise continuous functions characterized by a linear, spherical, and polynomial discontinuity interface identified by the zero-level set of given functions (see A.2 for details); specifically, we have 200 functions for each type. Then, we build the synthetic dataset of pairs $(\boldsymbol{g}^{\prime},\boldsymbol{p})$ from $\mathcal{G}_{2}$, running Algorithm 1 and detecting the zero-level sets containing the discontinuity interfaces of the functions in $\mathcal{G}_{2}$ (we use the detector $Z^{(150)}$ described in A.1). From now on, for the sake of simplicity, we will also refer to the discontinuity interfaces as “cuts”.
@@
  • Loss function: loss (17), with ${\mu_{0}}=0.5$ and ${\mu_{1}}=1.5$;
@@
Concerning the NN-based detectors, we use a leaky relu activation function [40], a Glorot Normal initialization for the weights [41], and (for the GINN-based detector) $F=15$ filters for the hidden layers; both MLPs and GINNs are characterized by layers of $N=|\mathcal{S}|$ units and by residual blocks (see [42]) for exploiting depth in the models, depth that is proportional to the diameter of the SGG. For more details about the model’s architectures, see A.4.
At the end of the training, we measure the loss and, for a better understanding, the Mean Absolute Error (MAE) of these models on the test set $\mathcal{P}_{2}$ (see Table 1). Looking at the values in the table, we see that the GINN’s prediction performances are slightly better than the MLP’s ones, at a reasonable cost in terms of trainable parameters. In particular, there are more trainable parameters in the GINN model because the GI hidden layers are endowed with $F=15$ filters; nonetheless, thanks to the sparsity of the adjacency matrix of the SGG, the number of trainable weights in the GINN is much smaller than $15$ times the number of MLP’s weights.
NN model # trainable param.s Loss function value MAE
MLP 52 9105291052\,91052 910 4.8678 0.0480
GINN 173 880173880173\,880173 880 3.5396 0.0393
@@
Now, we test the discontinuity detection abilities of these NN models if used as discontinuity detectors together with the sparse grid-based algorithm proposed (see Algorithm 2). In particular, we test them on four piece-wise continuous functions defined on $\Omega_{2}=[-1,1]^{2}$ (see Figure 7): $i$) a function with circular cut; $ii$) a function with Legendre polynomial cut; $iii$) a function with sinusoidal cut; $iv$) a function with one elliptic cut and two bow-shaped cuts. We point the attention of the reader to the fact that functions ($i$) and ($ii$) are of the same type of functions used to generate the synthetic dataset $\mathcal{D}_{2}$. On the other hand, functions ($iii$) and ($iv$) are functions of a different type with respect to the ones in $\mathcal{G}_{2}$; then, they are somehow “new” to the NN models and good performances on these functions witness good generalization abilities of the NNs. Moreover, looking at Figure 7, we can observe that test functions ($i$), ($ii$), and ($iv$) have discontinuity interfaces characterized by parts where the discontinuity jump is approximately zero, i.e., where the interface is not of (evident) discontinuity; then, they are important to understand how much sensitive the NN-based detectors are in identifying discontinuities.
@@
The performances of the NN-based detectors are measured according to a True Positive Rate (TPR). Let $T$ be the set of final troubled points detected by the method and let $\Lambda_{\min}$ be the minimum edge length used as stopping criterion (see Algorithms 1 and 2); then, the TPR with respect to each test functions is computed as follows:
@@
    let $\zeta$ be the zero-level set of the function characterizing the discontinuity interface of the test function. Let $\widetilde{\mathcal{S}}_{2}$ be the 2-dimensional sparse grid represented by the blue dots in Figure 8(a);
@@
For measuring the detection quality of our method, we do not consider a True Negative Rate (TNR) because a definition of true non-troubled point based on the distance from $\zeta$ (like true trouble points in TPR) would be misleading, even in the case of exact detectors. In fact, Algorithm 1 labels a point $\boldsymbol{x}_{i}$ as non-troubled point not because it is far from the discontinuity interface, but according to other criteria. Specifically, $\boldsymbol{x}_{i}$ is a non-troubled point either because no edge with $\boldsymbol{x}_{i}$ as vertex intersects the discontinuity interface (e.g., see the magenta triangle in Figure 8(b)) or because there exists an edge $e_{ij}=\{\boldsymbol{x}_{i},\boldsymbol{x}_{j}\}$, $\|{\boldsymbol{x}_{i}-\boldsymbol{x}_{j}}\|<\Lambda_{\min}$, such that $e_{ij}$ intersects the discontinuity interface but $\boldsymbol{x}_{j}$ is closer to the intersection than $\boldsymbol{x}_{i}$ (e.g., see orange the triangle in Figure 8(b)). In both cases, $\boldsymbol{x}_{i}$ is marked as non-troubled point not due to its distance from the discontinuity interface, but because there were better alternatives to it. To better understand these observations, we report in Figure 8(b) an example in which two true troubled points are detected; in the figure we highlight two examples of non-troubled points inside the area delimited by the grid $\widetilde{\mathcal{S}}_{2}$ in the middle of the picture: the orange triangle marks a non-troubled point that has an edge intersecting $\zeta$, and the magenta triangle denotes a non-troubled point without edges that intersect $\zeta$. If we would use these points as centers for the grid $\widetilde{\mathcal{S}}_{2}$ adopted to check the $\zeta$-intersection of the detected troubled points, the grid would intersect $\zeta$. For this reason a definition of true non-troubled point based on the distance from $\zeta$ like the true trouble points would be misleading.
@@
Figure 8: Left: SGG based on the sparse grid $\widetilde{\mathcal{S}}_{2}$. Right: use of the SGG for computing the TPR. Red segments are edges of the detector’s SGG; green circles are troubled points detected at the current step (dark green for final troubled points and light green for points used as grid centers in the next step); in the $\widetilde{\mathcal{S}}_{2}$ grid in the middle, orange and magenta triangles highlight two examples of non-troubled points sufficiently near to $\zeta$ (orange triangle if an edge intersects $\zeta$, magenta triangle if no edge intersects $\zeta$, see Remark 5.1).
We run Algorithm 2 on the four test functions using the NN-based detectors and using the same input parameters for all the cases; i.e., the same starting set $L$ of center and edge length pairs and the same $\Lambda_{\min}=2/2^{(h_{\max}+1)}=2^{-5}$ for stopping criterion ($h_{\max}=5$, maximum refinement level of the sparse grid). Moreover, since during the procedure some sparse grids may have points outside the domain $\Omega_{2}$, we modify Algorithm 2 in order to stop the search if a troubled point is detected outside the domain (i.e., it stops even if the stopping criterion with respect to $\Lambda_{\min}$ is not satisfied).
In Table 2, we report the TPRs, the number of troubled points detected, and the total number of points visited by the two NN-based discontinuity detectors, with respect to the test functions. Looking at the results reported in this table, we observe that both the NN-based detectors have extremely good TPRs (at least $98\%$). According to the test set performances (see Table 2), the GINN-based detector is more precise than the MLP-based one (higher TPRs and TPR equal to $100\%$ in three cases over four); moreover, more troubled points are detected. In particular, looking at the positioning of the detected troubled points (Figures 9 and 10), we can deduce that the higher sensitiveness of the GINN-based detector depends on its ability to detect troubled points even if the discontinuity jumps are very small.
@@
NN model Test function TPR Num. of troubled points Num. of visited points
@@ 5.1.1 Edge Detection Problem in Images
Black and white images can be modeled as functions $g:\Omega\subset\mathbb{R}^{2}\rightarrow[0,1]$, where $g(\boldsymbol{x})$ is the grayscale level of the pixel, for each pixel $\boldsymbol{x}\in\Omega$. Then, two-dimensional detectors can be used for detecting edges in pictures, since edges are discontinuity interfaces of $g$. Analogously to [3], we test our NN-based two-dimensional detectors on the edge detection problem of the Shepp-Logan phantom [43], varying the image resolution. In particular, let $M=(M_{ij})\in[0,1]^{r\times r}$ be the matrix representing the Shepp-Logan phantom image with resolution $r\times r$ pixels; then, we apply Algorithm 2 modeling the image as a function $g_{r}:\mathbb{R}^{2}\rightarrow[0,1]$ such that:
$g_{r}(\boldsymbol{x})=\begin{cases}M_{[x_{1}]+1,[x_{2}]+1}\,,\quad&\text{if }0\leq[x_{1}],[x_{2}]\leq r-1\\ 0\,,\quad&\text{otherwise}\end{cases}\,,$
where $[x]$ is the rounding to the nearest integer of $x$, for each $x\in\mathbb{R}$.
In Figure 11, we report the results obtained on $g_{r}$ using Algorithm 2 and the two NN-based discontinuity detectors, for each $r=512,1024$. Looking at these figures, we observe that the edge detection performances of the NN-based method increase with the resolution, similarly to the CNN detectors of [3]; however, we see that the GINN-based detector shows very good performances even in the resolution $512\times 512$ while the MLP-based detector obtains acceptable results (despite some imperfections) only for resolution $1024\times 1024$ (see Figure 11). Therefore, we have further experimental evidence of the better detection abilities of the GINN-based detector with respect to the MLP-based one.
A very interesting characteristic of our method applied to the edge detection problem can be observed in the rightmost column of Table 3, which reports the number of troubled points detected and the total number of points visited by the NN-based detectors, for each image resolution. Indeed, we observe that increments in the resolution do not imply a considerably higher number of points visited by the method. This characteristic is very important because it means that increments in the image resolution guarantee better results at almost the same cost; in particular, we have that the function evaluations done by the algorithm are (approximately) only $330^{2}$, that is much smaller than the number of pixels available ($512^{2}$ and $1024^{2}$, respectively).
(a) MLP, $512\times 512$
(b) GINN, $512\times 512$
(c) MLP, $1024\times 1024$
(d) GINN, $1024\times 1024$
@@
NN model Image Resolution Num. of troubled points Num. of visited points
@@ 5.2 Four-dimensional Discontinuity Detection
Following the considerations made before, in this subsection we focus on a GINN-based detector, as in general GINNs perform better than MLPs. The sparse grid we consider is $\mathcal{S}_{4}:=\mathcal{I}_{\rm sum}(8)$ in $\mathbb{R}^{4}$ ($N=401$ points). Then, we consider a randomly generated set $\mathcal{G}_{4}$ of 750 functions, again with linear, spherical, and polynomial cut (250 each). We use a larger set of functions (with respect to Section 5.1) because we use a less precise approximation of the exact detector for the creation of the synthetic dataset ($Z^{(3)}=\left.Z^{(t+1)}\right|_{t=2}$, see A.1 for details); indeed, we prefer a less precise detector to make the dataset creation process less expensive, even if at the cost of a less detailed dataset.
@@
NN model # trainable param.s Loss function value MAE
GINN 1 263 54012635401\,263\,5401 263 540 37.5907 0.0585
@@
Given the trained GINN-based four-dimensional detector, we test its abilities running Algorithm 2 ($\Lambda_{\min}=2/2^{(h_{\max})}=2^{-4}$) on a piece-wise continuous four-dimensional test function. Specifically, we use a test function $g_{\vartheta}:\Omega_{4}\subseteq\mathbb{R}^{4}\rightarrow\mathbb{R}$ with domain $\Omega_{4}=[-1,1]^{4}$, continuous pieces defined by four-dimensional Legendre polynomials (see (24), A.2), and as discontinuity interface a torus with size depending on $x_{4}$; i.e., the discontinuity interface of $g_{\vartheta}$ is characterized by the zero-level set of the function
@@
The results of Algorithm 2 for this four-dimensional case are very promising. In Table 5, we report the TPR, the number of troubled points detected, and the total number of points visited by the GINN-based detector; then, looking at this table, we see a TPR of $93.52\%$ for more than half a million of troubled points. Moreover, the total number of points evaluated during the algorithm execution is approximately equal to the number of points of a coarse regular grid of $60$ knots per dimension but the point distribution is more dense near to the discontinuity interface $\vartheta(\boldsymbol{x})=0$, instead of being equally distributed in $\Omega_{4}$.
In Figure 12, we report some examples of the troubled points identified by the GINN-based detector for the function $g_{\vartheta}$. In particular, for fixed values of $x_{4}$, we show cases where the detector identifies the discontinuity interface almost perfectly with (almost) no false troubled points (Figures 12(a) and 12(b)), with few true troubled points (Figure 12(c)), and with many false troubled points (Figure 12(d)).
NN model TPR Num. of troubled points Num. of visited points
@@
Figure 12: Visual examples of the troubled points detected by the GINN-based detector for $g_{\vartheta}$. Each picture is a visualization in $\mathbb{R}^{3}$ of $g_{\vartheta}$ for a fixed value of $x_{4}$. The colored surface is the size-changing torus $\vartheta(\boldsymbol{x})=0$ (see (21)). Black dots are true troubled points, red dots are false troubled points.